OpenStudy (anonymous):

The volume of a hexagonal prism is 12 sqr root 3 in3. What is the height of the prism if the base area is 3 sqr root 3 in3?

OpenStudy (anonymous):

@TuringTest

OpenStudy (eyust707):

The volume of any "prism" is: \[(Area \space of \space Base)*(Height)\]

OpenStudy (anonymous):

so what do I plug in?

OpenStudy (eyust707):

\[Volume = (Area \space of \space Base)*(Height)\] Volume = \(12\sqrt{3} \) Area of base = \(3\sqrt{3}\)

OpenStudy (eyust707):

Solve for Height

OpenStudy (anonymous):

6 sqr root 3?

OpenStudy (eyust707):

you sure...

OpenStudy (anonymous):

no... how do I solve for it?

OpenStudy (eyust707):

Divide both sides by \(3 \sqrt{3}\) \[12\sqrt{3} \over 3\sqrt{3}\]

OpenStudy (eyust707):



OpenStudy (eyust707):